OpenStudy (anonymous):

Please Help! 8th grade math Line Q is represented by the following equation: 2x + y = 11 Which equation completes the system that is satisfied by the solution (3, 5)?

OpenStudy (anonymous):

A. x + 2y = 15 B. x + y = 15 C. x + y = 8 D. x − y = 2

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

@UsukiDoll @skullpatrol